OpenStudy (anonymous):

What is the slope of the line tangent to the curve y+2 = (x^2/2) - 2siny at the point (2,0)?

OpenStudy (australopithecus):

y + 2 = (x^(2)/2) rearange the equation take the derivative

OpenStudy (anonymous):

i know the derivative is teh slope of the line to a tangent but m not sure how to differentiate that

OpenStudy (anonymous):

u forgot the -2siny

OpenStudy (australopithecus):

sorry f'(x)(x-a) + f(x) = m

OpenStudy (anonymous):

i know the equation but just not how 2 differentiate f(x)....tahts wat i need help with

OpenStudy (australopithecus):

alright, y+2 = (x^2/2) - 2sin(y) so we are taking the derivative y in respect to x so we have dy/dx use chain rule on y so y' = 2x/2 - 2cos(y)*y'

OpenStudy (anonymous):

the 2 disapears cuz the derivative of a constant is 0, right?

OpenStudy (australopithecus):

Yes to your question Now rearrange it to solve for y' y' = 2x/2 - 2cos(y)*y' 0 = x - 2cos(y)y' - y' - x = 2cos(y)y' - y' -x = y'(2cos(y) - 1) -x/(2cos(y) - 1) = y'

OpenStudy (australopithecus):

we know when f(2) = 0 so thus y = 0 so when f'(2) = -2/(2cos(0)-1)

OpenStudy (anonymous):

what happened to the 2x/2 after u made it equal to 0? why did it cancel?

OpenStudy (australopithecus):

2/2 = 1

OpenStudy (australopithecus):

x(2/2) = x

OpenStudy (anonymous):

OH righ? sorry i see it now

OpenStudy (anonymous):

when i plug in 0 i get -2/0 = 0

OpenStudy (australopithecus):

you cant divide by 0

OpenStudy (anonymous):

so what do i do?

OpenStudy (australopithecus):

f'(2) = -2/(2cos(0)-1) cos(0) = 1 thus f'(2) = -2/(2(1)-1) = -2/-1 = 2 f'(2) = 2

OpenStudy (anonymous):

how did u get cos(0) =1

OpenStudy (australopithecus):

I have the unit circle memorized

OpenStudy (australopithecus):

 cos = x sin = y

OpenStudy (anonymous):

i have the asnwer key 2 the question but its says the asnwer is 2/3...its just taht idk how to get that answer

OpenStudy (australopithecus):

I made a mistake and didnt carry the negative f'(2) = -2/(-2cos(0)-1) f'(2) = 2/3

OpenStudy (australopithecus):

that is your answer

OpenStudy (australopithecus):

y' = 2x/2 - 2cos(y)*y' 0 = x - 2cos(y)y' - y' - x = - 2cos(y)y' - y' - x = y'(- 2cos(y) - 1) - x/(- 2cos(y) - 1) = y'

OpenStudy (anonymous):

YOU ARE AWESOME! I TOTALLY UNDERSTOOD THE WHOLE PROCESS. Sorry 4 askingso many questions. Sometimes i dont see things right away

OpenStudy (australopithecus):

its fine engagement is a good thing